OpenStudy (anonymous):

How do I know if the foci should be in the x axis or y axis? This is about graphing ellipses.

OpenStudy (btaylor):

The foci is on the major axis. An ellipse has the form \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\]Whichever is bigger, a or b, determines which is the major axis. If a>b, then it is on the x-axis. If a<b, it is on the y-axis. Does that make sense?

OpenStudy (anonymous):

Thank you so much! @BTaylor

OpenStudy (btaylor):

you're welcome